# Predicting function of evolutionarily implausible DNA sequences

### v1

@@ Appendix C Model Performance Evaluation
We systematically evaluated the impact of tokenization strategies and model sizes on the performance of gLMs. As shown in Appendix Figure 5, we benchmarked the PDLLM model series (Liu et al., 2025a), which differ only in their tokenization choices. Our analysis revealed that the choice of tokenization profoundly influences model success rates: byte pair encoding (BPE) shows a stable performance across models. In contrast, single-token schemes performed poorly, while the 6-mer representation exhibited intermediate performance, with model-specific variability. To assess the effect of model size, we compared models that vary only in parameter size. Notably, we found no consistent relationship between model size and task performance: in several cases, smaller models performed on par with, or even outperformed, their larger counterparts.
Figure 5: Comparative evaluation of gLMs across tokenization methods (left) and model sizes (right).
Furthermore, while representative models were selected for the main analysis, we conducted an extensive evaluation of various model versions across all datasets. As shown in Appendix Figure 6, the Evo series models (Nguyen et al., 2024; Merchant et al., 2024; Brixi et al., 2025) consistently outperformed all others in the Nullsettes tasks, with nearly every Evo variant ranking at the top. This advantage likely reflects the superior architectural design and scaling strategy employed by the Evo models. Among non-Evo models, METAGENE-1 (Liu et al., 2025b) achieved the highest performance, likely due to its training on over 1.5 trillion base pairs of DNA and RNA sequences derived from wastewater metagenomes. The unique breadth and diversity of metagenomic data may provide this model with an enhanced ability to generalize across sequence contexts. The Nucleotide Transformer (NT) (Dalla-Torre et al., 2024) series followed, ranking third overall. These findings highlight the impact of both model architecture and training data diversity on generalization performance across regulatory sequence prediction tasks.
@@
gLM2 (Cornman et al., 2024) OMG: encodes genomic scaffolds with both amino-acid (CDS) and DNA tokens (315B tokens). MLM Char-level tokens (DNA: lowercase, AA: uppercase) Transformer, 650M params Pretrained with a 4096 token context window

### v2

@@ Appendix C Model Performance Evaluation
While representative models were selected for the main analysis, we conducted an extensive evaluation of various model versions across all datasets. As shown in Appendix Figure 5, the Evo series models (Nguyen et al., 2024; Merchant et al., 2024; Brixi et al., 2025) consistently outperformed all others in the Nullsettes tasks, with all the Evo variants ranking at the top. This advantage likely reflects the superior architectural design and scaling strategy employed by the Evo models. Among non-Evo models, METAGENE-1 (Liu et al., 2025b) achieved the highest performance, likely due to its training on over 1.5 trillion base pairs of DNA and RNA sequences derived from wastewater metagenomes. The unique breadth and diversity of metagenomic data may provide this model with an enhanced ability to generalize across sequence contexts. The Nucleotide Transformer (NT) (Dalla-Torre et al., 2024) series followed, ranking third overall. These findings highlight the impact of both model architecture and training data diversity on generalization performance across regulatory sequence prediction tasks.